OpenStudy (anonymous):

Help please I'll give medalsssss ~*$(%(#@* Max is in a control tower at a small airport. He is located 50 feet above the ground when he spots a small plane on the runway at an angle of depression of 27. What is the distance of the plane from the base of the tower? Round to the nearest foot. A. 25 feet B. 110 feet C. 56 feet D. 98 feet It's not A, because that was my first choice but I got it wrong and I just want to know why and what the actual answer is. Thank you!

OpenStudy (anonymous):

why did you choose A? how did you get that answer?

OpenStudy (anonymous):



OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

tan(27) = x/50 => x = 50 * tan(27) => x = 50*0.5097 => x = 25.47ft = 25ft.

OpenStudy (anonymous):

\[\tan(63)=\frac{x}{50}\] so \[x=50\tan(63)\]

OpenStudy (anonymous):

yeah, but the angle of depression means what i wrote in the picture above

OpenStudy (anonymous):

I don't have a calculator so that's been giving me a bit of trouble

OpenStudy (anonymous):

i googled it look carefully, it is not \(\tan(27) \) it should be \(\tan(63)\)

OpenStudy (anonymous):

of course you have a calculator, you are on a computer !! http://www.wolframalpha.com/input/?i=50tan%2863%29

OpenStudy (anonymous):

hahahahaha true

OpenStudy (anonymous):

looks like about 98

OpenStudy (anonymous):

You're wonderful. Thank you for the help!

OpenStudy (anonymous):

yw hope it is clear why you use 63 and not 27

OpenStudy (anonymous):

Can you help me with some others? I'll post new questions so that I can reward more medals if you wish.

OpenStudy (anonymous):

i have enough medals, if you want to ask one here go ahead

OpenStudy (anonymous):

Okay, I have to review my assignments first. (: Thank youuu

OpenStudy (anonymous):

oh ok, then probably best to post in a new thread

OpenStudy (anonymous):

thnx a lot cause i had gotten a too and had gotten it wrong thnx